# Cryptomite: A versatile and user-friendly library of randomness extractors

### v1

@@ 3.2.5 Composing extractors
Extractors can be composed together to form new extractors with improved parameters. One can append a strong seeded extractor to a strong two-source extractor, taking the two-source extractor output as a seed and the two-source weak seed as the extractor input. Doing so, one can construct a $(k_{1},k_{2},\epsilon_{2}+\epsilon_{s},n_{1},n_{2},m+\max\{k_{1},k_{2}\})$ two-source extractor, which significantly increases the output length. For more details, see Lemma 38 in [11] and the discussion in Section 4.6 of [16].
@@ 3.3 Our New $\mathsf{Circulant}$ Construction
We now expose the details of our new extractor construction $\mathsf{Circulant}$.
@@
Let $x=x_{0},\ldots,x_{n-2}\in\{0,1\}^{n-1}$ and $y=y_{0},\ldots,y_{n-1}\in\{0,1\}^{n}$ where $n$ is prime. The function $\mathsf{Circulant}(x,y):\{0,1\}^{n-1}\times\{0,1\}^{n}\rightarrow\{0,1\}^{m}$ is implemented by:
@@
    Set $x^{\prime}=[x,0]\in\{0,1\}^{n}$, where [,][,][ , ] denotes concatenation (i.e. $x^{\prime}_{n-1}=0$).
@@
    where the matrix-vector multiplication $\mathrm{circ}(x^{\prime})y$ is taken mod 2 (in each component of the resulting vector) and the subscript ${0:m-1}$ denotes the first $m$ elements of the vector. The term $\mathrm{circ}(x^{\prime})$ is the $n\times n$ circulant matrix generated by $x^{\prime}=x^{\prime}_{0},\ldots x^{\prime}_{n-1}$,
@@
The function $\mathsf{Circulant}(X,Y):\{0,1\}^{n-1}\times\{0,1\}^{n}\rightarrow\{0,1\}^{m}$ in Definition 9 is a strong in $Y$, classical-proof and quantum-proof $(k_{1},k_{2}=n,\epsilon,n-1,n,m)$-seeded extractor for prime $n$, with output length
@@
Using Theorem 1 to build a two-source extension that is (classical-proof and) quantum-proof in the product sources model, we get that our $\mathsf{Circulant}$ extractor has an output length
@@
and, applying Theorem 4,
$m\leq\frac{1}{5}\left(k_{1}+k_{2}-n+4\log_{2}(\frac{4}{3})\right)$ (24)
in the Markov model.
@@
The exact $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ extractor [15] $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}(x,y):\{0,1\}^{n}\times\{0,1\}^{n}\to\{0,1\}^{m}$ construction is given in Section A.3, whilst we give its idea here. The extractor input $x=x_{0},x_{1},\ldots,x_{n-1}\in\{0,1\}^{n}$, and the (weak) seed $y=y_{0},y_{1},\ldots,y_{n-1}\in\{0,1\}^{n}$ must both be of equal length $n$. The extractor construction uses a set of $m$ $n\times n$ matrices, which we label $A_{0},...,A_{m-1}$, with entries in $\{0,1\}$ (i.e. bits), which must be chosen such that, for any subset $B\subseteq A_{0},...,A_{m-1}$, the sum of the matrices in the subset $B$ has rank at-least $n-r$ for some small constant $r$ (which acts as a penalty to the output length). The extractor output of $m$ bits is then
$\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}(x,y)=[(A_{0}x)\cdot y,(A_{2}x)\cdot y,...,(A_{m-1}x)\cdot y]$ (25)
where $\cdot$ is the inner product modulo 2. We give such a set of matrices, based on “Cyclic Shift matrices” from [15] Section 3.2, and a version of Equation 25 that can be implemented in $O(n\log(n))$ computation time, based on [16] in Section A.3. This construction requires that $x\neq\{0\}^{n}$ or $\{1\}^{n}$ and that $n$ is a prime with 2 as a primitive root [^9] A list of all primes with 2 as a primitive root up to $10^{6}$ can be found at https://github.com/CQCL/cryptomite/blob/main/na_set.txt. We also provide a function which allows the user to compute the closest prime with 2 as primitive root, as well as individually the closest above and below the input length, see the utility functions https://github.com/CQCL/cryptomite/blob/main/cryptomite/utils.py.. For the input lengths considered in this work (i.e. up to $10^{12}$), this set is sufficiently dense in the natural numbers for all practical purposes. We also note that fixing the input length is a pre-computation that can be done very efficiently.
@@
Our implementation of the $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ extractor is a strong classical-proof $(k_{1},n,\epsilon,n,n,m)$-seeded randomness extractor with
@@
and a classical-proof $(k_{1},k_{2},\epsilon,n,n,m)$-two-source randomness extractor, strong in the second source (with $k_{2}$), with
@@
The $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ extractor becomes a strong quantum-proof (in the Markov model) $(k_{1},n,\epsilon,n,n,m)$-seeded randomness extractor, where
$m\leq\frac{1}{5}\left(k_{1}+8\log_{2}\epsilon+4\log_{2}(\frac{4}{3})+1\right),$ (28)
and a strong quantum-proof (in the Markov model) $(k_{1},k_{2},\epsilon,n,n,m)$-two-source randomness extractor, where
$m\leq\frac{1}{5}\left(k_{1}+k_{2}-n+8\log_{2}\epsilon+4\log_{2}(\frac{4}{3})+1\right).$ (29)
@@ 3.5 Toeplitz
$\mathrm{toep}(y)=\begin{bmatrix}y_{0}&y_{1}&\ldots&y_{n-1}\\ y_{n+1}&y_{0}&\ldots&y_{n-2}\\ \vdots&&\ddots&\vdots\\ y_{n+m-3}&y_{n+m-4}&\ldots&y_{n-m+1}\\ y_{n+m-2}&y_{n+m-3}&\ldots&y_{n-m}\end{bmatrix}$ (36)
is the $n\times m$ Toeplitz matrix generated from $y$. The full implementation details, in complexity $O(n\log(n))$, is given in Section A.4. As it was shown to be directly quantum-proof [49, 17], we state the claims as a single theorem.
@@
The $\mathsf{Toeplitz}$ extractor is a strong classical- and quantum-proof $(k_{1},n+m-1,\epsilon,n,n+m-1,m)$-seeded randomness extractor, where
@@
Using the two weak sources extension (Theorem 1), a classical- and quantum-proof $(k_{1},k_{2},\epsilon,n,n+m-1,m)$-two-source randomness extractor strong in the second source (with $k_{2}$), where
@@
in the product sources model and, using Theorem 4,
$m\leq\frac{1}{6}\left(k_{1}+k_{2}-n+8\log_{2}\epsilon+4\log_{2}(\frac{4}{3})+1\right),$ (39)
in the Markov model. Note that $k_{2}$ may be bigger than $n$, due to the fact that the length of the (weak) seed is $n+m-1$.
Modified Toeplitz Extractor – To reduce the seed length, one can use modified Toeplitz matrices (Subsection B-B in [8]), which allow a reduction from $n+m-1$ to $n-1$. The authors also prove the security of the two-source extension against quantum side-information on the weak input only (which we reproduced as Theorem 1) and in [36] the security is extended against product quantum side information (i.e. side-information on both inputs). One could further extend it to allow for more general forms of quantum side information (in the Markov model) using Theorem 4.
@@ 3.6 Trevisan
The $\mathsf{Trevisan}$ extractor [7] $\mathsf{Trevisan}(x,y):\{0,1\}^{n}\times\{0,1\}^{d}\to\{0,1\}^{m}$ is important because of its asymptotic logarithmic seed length $d=O(\log(n))$. At a high level, it is built from two components: (1) a weak design which expands the uniform seed into several separate bit strings (chunks) that have limited overlap, in the sense that only a subset of bits from any two strings match, and (2) a 1-bit extractor which combines each of the chunks, in turn, with the weak random source to produce one (near-)perfect bit. By applying the 1-bit extractor to each expanded chunk, the Trevisan extractor produces multiple bits of near-perfect randomness. Its drawback is its computation time, which often prohibits its use in practice. The implementation, based on the building blocks of [19], in complexity $O(n^{2}\log(\log(\frac{n^{2}}{\epsilon}))$, can be found in Section A.5 and has a seed size, $d$, of
@@
The $\mathsf{Trevisan}$ extractor is a strong (classical-proof or) quantum-proof $(k_{1},d,\epsilon,n,d,m)$-seeded randomness extractor, with
@@
Using the two-weak-source extension (Theorem 1), the $\mathsf{Trevisan}$ extractor is a classical-proof $(k_{1},k_{2},\epsilon,n,d,m)$-two-source randomness extractor strong in the second source (with $k_{2}$), where
@@
and, using Theorem 4, a quantum-proof
$m\leq\frac{1}{10}\left(k_{1}+4(k_{2}-d)-4\log_{2}(m)+18\log_{2}\epsilon+9\log_{2}(\frac{4}{3})-6\right),$ (45)
two-source extractor in the Markov model.
@@ 3.7 Von Neumann
$\Pr(X_{2i}=0|E)=\Pr(X_{2i+1}=0|E)=p_{i}\hskip 28.45274pt\forall i\in[0,\ldots,\Big{\lfloor}\frac{n}{2}\Big{\rfloor}-1]$ (46)
for some $p_{i}\in[0,1]$ and some adversary side information $E$ (or alternatively, for a quantum-proof extractor, $\mathcal{E}$). Let $j=0,\ldots,\lfloor n/2\rfloor-1$, the $j$th output bit of the $\mathsf{Von\hskip 2.84544ptNeumann}$ extractor is then given by
$\mathsf{Von\hskip 2.84544ptNeumann}(\mathbf{x})_{j}=\begin{cases}x_{2j}&\quad\text{if}\hskip 14.22636ptx_{2j}\neq x_{2j+1}\\ \emptyset&\quad\text{otherwise},\end{cases}$ (47)
where the empty set $\emptyset$ denotes that there is no bit output in that round. Note that, contrary to the other constructions, the output size of the $\mathsf{Von\hskip 2.84544ptNeumann}$ extractor is probabilistic. For example, a valid input $X$ is one that is IID with an unknown bias $p_{i}=p$ for all $i$. In this case, the approximate (due to finite sampling from $X$) output length is given by
@@
which gives an entropy loss $k-m$ of $(1+(1-p)p/\log_{2}(\max\{p,1-p\}))k_{1}\geq 3k_{1}/10$ i.e. implying at-least $30\%$ entropy loss. This is significantly worse than the entropy loss when using the same input to an optimal seeded extractor, where $m=\lfloor-n\log_{2}(\max\{p,1-p\})-2\log_{2}(\frac{1}{\epsilon})\rfloor$ i.e. approximately $0\%$ entropy loss when $\epsilon$ is constant. Other extractors based on this construction, but optimising the output size by recycling unused bits, allow for this entropy loss to be improved, including the Elias and Peres extractor [51] and the generalised Von Neumann extractor [52]. However, they are still unable to achieve the same entropy loss that is possible with seeded extractors, in general. For completeness, we provide a pseudo-code implementation in Section A.6 with complexity $O(n)$.
@@ 4.1 Cryptomite  for Privacy Amplification in QKD
In QKD, the goal is for two parties, Alice and Bob, to generate a shared secret key that is unpredictable to an adversary. After rounds of quantum and classical communication, Alice and Bob share an identical shared raw key which is only partially secret to the adversary (i.e. after state sharing and measurement, sifting, parameter estimation and error reconciliation). Randomness extractors are used for privacy amplification, a subroutine which transforms the raw key (that has some min-entropy) into a final secret key that is (almost) completely secret to adversary (i.e. $\epsilon$-perfectly random to the adversary). In standard protocols, this task is performed using a strong seeded extractor, which requires both Alice and Bob to have a shared (near)-perfect seed.
We consider the continuous variable QKD demonstration of [53], which aims at security against quantum adversaries and therefore requires a quantum-proof extractor. From the experiment, Alice and Bob obtain $n=1.738\times 10^{9}$ bits of shared raw key as input for privacy amplification [^10] Since the extractor input length is bigger than $2^{29}$, the extractor implementation will use the bigNTT, implying a throughput of 34×3-4$\times$3 - 4 × less than that shown in Section 2.3.. After evaluating the total min-entropy of the raw key, the authors compute the final secret key length of $m=41378264$ bits based on their given security parameters (including an extraction error of $\epsilon=10^{-10}$) when using their $\mathsf{Toeplitz}$ extractor implementation. This extraction requires a perfect seed of length $d=m+n-1=1.738\times 10^{9}+41378263$ and can be performed in a few lines of Python code using Cryptomite, as shown in Figure 5.
Extensions – The $\mathsf{Circulant}$ extractor can be used to generate the same amount of shared secret key, whilst substantially reducing the size of the seed. The code to perform this is given in Figure 6. The setup can be further improved, as in the experiment a quantum-RNG (QRNG) is used to generate the extractor seed [55] and its quality relies on the correct characterisation and modelling of the components in the device (an assumption which was shown to be potentially problematic in [56]). We show how to relax this assumption by assuming that the min-entropy rate of the QRNG output is only $r=0.99$, i.e. $k_{2}=r\cdot n_{2}$, and adjusting the output length of the extractor accordingly (in the quantum-proof product sources model), see Figure 3. The code to perform the parameter calculation is given in Figure 6, but it could also easily be performed using the calculation module from_params (see documentation).
@@
%**** main.tex Line 800 ****(after sifting, error correction and parameter estimation).
@@
# To use Circulant, the seed length needs to be a prime, we
# find the one larger than or equal to n_1 + 1 (see Def. 9):
n_adjusted = cryptomite.utils.next_prime(n_1 + 1)
@@
%**** main.tex Line 825 ****k_2 = r * n_adjusted
# Compute the extractors output length (product sources

### v3

@@ 3.2.5 Composing extractors
Extractors can be composed together to form new extractors with improved parameters. One can append a strong seeded extractor to a strong two-source extractor, taking the two-source extractor output as a seed and the two-source weak seed as the extractor input. Doing so, one can construct a $(n_{1},k_{1},n_{2},k_{2},m+\max\{k_{1},k_{2}\},\epsilon_{2}+\epsilon_{s})$-two-source extractor, which significantly increases the output length. For more details, see Lemma 38 in [13] and the discussion in Section 4.6 of [18].
4 Library in detail
In this section, we concretely define the extractors in Cryptomite, providing implementation details and parameter calculations.
@@ 4.1 Our new $\mathsf{Circulant}$ construction
We now expose the details of our new extractor construction, $\mathsf{Circulant}$.
@@
Let $x=x_{0},\ldots,x_{n-2}\in\{0,1\}^{n-1}$ and $y=y_{0},\ldots,y_{n-1}\in\{0,1\}^{n}$ where $n$ is prime with primitive root 2.101010A list of all primes with 2 as a primitive root up to $10^{6}$ can be found at https://github.com/CQCL/cryptomite/blob/main/na_set.txt. We also provide a function which allows the user to compute the closest prime with 2 as primitive root, as well as individually the closest above and below the input length, see the utility functions https://github.com/CQCL/cryptomite/blob/main/cryptomite/utils.py. The function $\mathsf{Circulant}(x,y):\{0,1\}^{n-1}\times\{0,1\}^{n}\rightarrow\{0,1\}^{m}$ is implemented by:
@@
    Set $x^{\prime}=[x,0]\in\{0,1\}^{n}$, where $[,]$ denotes concatenation (i.e., $x^{\prime}_{n-1}=0$).
@@
    where the matrix-vector multiplication $\mathrm{circ}(x^{\prime})y$ is taken mod 2 (in each component of the resulting vector) and the subscript 0:m1{0:m-1} denotes the first $m$ elements of the vector. The term $\mathrm{circ}(x^{\prime})$ is the $n\times n$ circulant matrix generated by $x^{\prime}=x^{\prime}_{0},\ldots x^{\prime}_{n-1}$,
@@
The function $\mathsf{Circulant}(X,Y):\{0,1\}^{n-1}\times\{0,1\}^{n}\rightarrow\{0,1\}^{m}$ in Definition 9 is a strong in $Y$, classical-proof and quantum-proof $(n_{1}=n-1,k_{1},n_{2}=n,k_{2}=n,m,\epsilon)$-seeded extractor for prime with primitive root 2 $n$, with output length
@@
Using Theorem 1 to build a two-source extension that is (classical-proof and) quantum-proof in the product source model, our $\mathsf{Circulant}$ extractor has an output length
@@
and, by Corollary 1, quantum-proof in the Markov model with the same output length.
@@
The exact $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ extractor [17] $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}(x,y):\{0,1\}^{n}\times\{0,1\}^{n}\to\{0,1\}^{m}$ construction is given in Section A.4, whilst we give its idea here. The extractor input $x=x_{0},x_{1},\ldots,x_{n-1}\in\{0,1\}^{n}$, and the (weak) seed $y=y_{0},y_{1},\ldots,y_{n-1}\in\{0,1\}^{n}$ must both be of equal length $n$. The extractor construction uses a set of $m$ $n\times n$ matrices, which we label $A_{0},...,A_{m-1}$, with entries in $\{0,1\}$ (i.e. bits), which must be chosen such that, for any subset $B\subseteq A_{0},...,A_{m-1}$, the sum of the matrices in the subset $B$ has rank at least nrn-r for some small constant $r$ (which acts as a penalty to the output length). The extractor output of $m$ bits is then
$\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}(x,y)=[(A_{0}x)\cdot y,(A_{1}x)\cdot y,...,(A_{m-1}x)\cdot y]$ (27)
where $\cdot$ is the inner product modulo 2. We give such a set of matrices, based on “Cyclic Shift matrices” from [17] Section 3.2, and a version of Equation 27 that can be implemented in $O(n\log(n))$ computation time, based on [18] in Section A.4. This construction requires that $x\neq\{0\}^{n}$ or $\{1\}^{n}$ and that $n$ is a prime with 2 as a primitive root. For the input lengths considered in this work (i.e. up to $10^{12}$), this set is sufficiently dense in the natural numbers for all practical purposes. We also note that fixing the input length is a pre-computation that can be done very efficiently.
@@
Our implementation of the $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ extractor is a strong classical-proof $(n_{1}=n,k_{1},n_{2}=n,k_{2}=n,m,\epsilon)$-seeded randomness extractor with
@@
and a classical-proof $(n_{1}=n,k_{1},n_{2}=n,k_{2},m,\epsilon)$-two-source randomness extractor, strong in the weak seed (with min-entropy $k_{2}$), with
@@
The $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ extractor becomes a strong quantum-proof (in the Markov model) $(n_{1}=n,k_{1},n_{2}=n,k_{2}=n,m,\epsilon)$-seeded randomness extractor, with
$m\leq\frac{1}{5}\left(k_{1}+8\log_{2}\epsilon+4\log_{2}\left(\frac{4}{3}\right)+1\right),$ (30)
and a strong quantum-proof (in the Markov model) $(n_{1}=n,k_{1},n_{2}=n,k_{2},m,\epsilon)$-two-source randomness extractor, where
$m\leq\frac{1}{5}\left(k_{1}+k_{2}-n+8\log_{2}\epsilon+4\log_{2}\left(\frac{4}{3}\right)+1\right).$ (31)
@@ 4.3 Toeplitz
$\mathrm{toep}(y)=\begin{bmatrix}y_{0}&y_{n+m-2}&\ldots&\ldots&y_{m}\\ y_{1}&y_{0}&\ddots&\ddots&\vdots\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ y_{m-1}&y_{m-2}&\ldots&\ldots&.\end{bmatrix}$ (33)
is the $n\times m$ Toeplitz matrix generated from $y$. The full implementation details, in computation time $O(n\log(n))$, is given in Section A.5. As it was shown to be directly quantum-proof [60, 19], we state the claims as a single theorem.
@@
The $\mathsf{Toeplitz}$ extractor is a strong classical- and quantum-proof $(n_{1}=n,k_{1},n_{2}=n+m-1,k_{2}=n+m-1,m,\epsilon)$-seeded randomness extractor, where
@@
Using the two weak sources extension (Theorem 1), a classical- and quantum-proof $(n_{1}=n,k_{1},n_{2}=n+m-1,k_{2},m,\epsilon)$-two-source randomness extractor strong in the weak seed (with min-entropy $k_{2}$), where
@@
in the product source model and also in the Markov model (by Corollary 1). Note that $k_{2}$ may be bigger than $n$, due to the fact that the length of the (weak) seed is n+m1n+m-1.
Modified Toeplitz Extractor – To reduce the seed length, one can use modified Toeplitz matrices (Subsection B-B in [7]), which allow a reduction from n+m1n+m-1 to n1n-1. The authors also prove the security of the two-source extension against quantum side information on the weak input only (which we reproduced as Theorem 1) and in [46] the security is extended against product quantum side information (i.e. side information on both inputs). One could further extend it to allow for more general forms of quantum side information (in the Markov model) using Theorem 4.
@@ 4.4 Trevisan
The $\mathsf{Trevisan}$ extractor [8] $\mathsf{Trevisan}(x,y):\{0,1\}^{n}\times\{0,1\}^{d}\to\{0,1\}^{m}$ is important because of its asymptotic logarithmic seed length $d=O(\log(n))$. At a high level, it is built from two components: (1) a weak design which expands the uniform seed into several separate bit strings (chunks) that have limited overlap, in the sense that only a subset of bits from any two strings match, and (2) a 1-bit extractor which combines each of the chunks, in turn, with the weak input to produce one (near-)perfect bit. By applying the 1-bit extractor to each expanded chunk, the Trevisan extractor produces multiple bits of near-perfect randomness. Its drawback is its computation time, which often prohibits its use in practice. The implementation, based on the building blocks of [21], in computation time $O(n^{2}\log(\log(\frac{n^{2}}{\epsilon})))$, can be found in Section A.6 and has a seed length, $d$, of
@@
The $\mathsf{Trevisan}$ extractor is a strong (classical-proof or) quantum-proof $(n_{1}=n,k_{1},n_{2}=d,k_{2}=d,m,\epsilon)$-seeded randomness extractor, with
@@
Using the two-weak-source extension (Theorem 1), the $\mathsf{Trevisan}$ extractor is a classical-proof $(n_{1}=n,k_{1},n_{2}=d,k_{2},m,\epsilon)$-two-source randomness extractor strong in the weak seed (with min-entropy $k_{2}$), where
@@
and, using Theorem 4, a quantum-proof $(n_{1}=n,k_{1},n_{2}=d,k_{2},m,\epsilon)$-two-source randomness extractor in the Markov model with
$m\leq\frac{1}{10}\left(k_{1}+4(k_{2}-d)-4\log_{2}(m)+18\log_{2}\epsilon+9\log_{2}\left(\frac{4}{3}\right)-6\right).$ (41)
@@ 4.5 Von Neumann
$\Pr(X_{2i}=0|E)=\Pr(X_{2i+1}=0|E)=p_{i},$ (42)
for all $i\in\{0,\ldots,\lfloor n/2\rfloor-1\}$, some $p_{i}\in[0,1]$ and some adversary side information $E$ (or alternatively, for a quantum-proof extractor, side information $\mathcal{E}$). Let $j=0,\ldots,\lfloor n/2\rfloor-1$, the $j$th output bit of the $\mathsf{Von\hskip 2.84544ptNeumann}$ extractor is then given by
$\mathsf{Von\hskip 2.84544ptNeumann}(\mathbf{x})_{j}=\begin{cases}x_{2j}&\quad\text{if}\hskip 14.22636ptx_{2j}\neq x_{2j+1}\\ \emptyset&\quad\text{otherwise}\end{cases},$ (43)
where the empty set $\emptyset$ denotes that there is no output bit in that round. Note that, contrary to the other constructions, the output length of the $\mathsf{Von\hskip 2.84544ptNeumann}$ extractor is probabilistic. For example, a valid input $X$ is one that is IID with an unknown bias $p_{i}=p$ for all $i$. In this case, the approximate (due to finite sampling from $X$) output length is given by
@@
which gives an entropy loss kmk-m of $(1+(1-p)p/\log_{2}(\max\{p,1-p\}))k_{1}\geq 3k_{1}/10$ i.e. implying at least $30\%$ entropy loss. This is significantly worse than the entropy loss when using the same input to an optimal seeded extractor, where $m=\lfloor-n\log_{2}(\max\{p,1-p\})-2\log_{2}(\frac{1}{\epsilon})\rfloor$ i.e. approximately $0\%$ entropy loss when $\epsilon$ is constant. Other extractors based on this construction, but optimising the output length by recycling unused bits, allow for this entropy loss to be improved, including the Elias and Peres extractor [62] and the generalised Von Neumann extractor [63]. However, they are still unable to achieve the same entropy loss that is possible with seeded extractors, in general. For completeness, we provide a pseudo-code implementation in Section A.7 with computation time $O(n)$.
@@ 5 Cryptomite: Examples with code
Remark that all the examples that we give in this section also showcase the limitation of the Trevisan extractor (and others) in practice. Indeed, for these examples the input lengths are too long for implementations that do not have quasi-linear computation time $O(n\log(n))$, i.e. that the usual notion of polynomial efficiency is irrelevant in practical quantum cryptography. We refer the reader to Figure 4 for a numerical example of the performance with different input lengths.
@@ 5.1 Cryptomite  for privacy amplification in QKD
In QKD, the goal is for two parties, Alice and Bob, to generate a shared secret key that is unpredictable to an adversary. After rounds of quantum and classical communication, Alice and Bob share an identical shared raw key which is only partially secret to the adversary (i.e. after state sharing and measurement, sifting, parameter estimation and error reconciliation). Randomness extractors are used for privacy amplification, a subroutine which transforms the raw key (that has some min-entropy) into a final secret key that is (almost) completely secret to adversary (i.e. $\epsilon$-perfectly random to the adversary). In standard protocols, this task is performed using a strong seeded extractor, which requires both Alice and Bob to have a shared (near-)perfect seed.
We consider the continuous-variable QKD demonstration of [64], which aims at security against quantum adversaries and therefore requires a quantum-proof extractor. From the experiment, Alice and Bob obtain $n=1.738\times 10^{9}$ bits of shared raw key as input for privacy amplification [^11] Since the extractor input length is bigger than $2^{29}$, the extractor implementation will use the bigNTT, implying a throughput of $3-4\times$ less than that shown in Section 2.3.. After evaluating the total min-entropy of the raw key, the authors compute the final secret key length of $m=41378264$ bits based on their given security parameters (including an extraction error of $\epsilon=10^{-10}$) when using their $\mathsf{Toeplitz}$ extractor implementation. This extraction requires a perfect seed of length $d=m+n-1=1.738\times 10^{9}+41378263$ and can be performed in a few lines of Python code using Cryptomite, as shown in Figure 5.
Extensions – The $\mathsf{Circulant}$ extractor can be used to generate the same amount of shared secret key, whilst substantially reducing the length of the seed. The code for this is given in Figure 6. The setup can be further improved, as in the experiment a quantum-RNG (QRNG) is used to generate the extractor seed [66] and its quality relies on the correct characterisation and modelling of the components in the device (an assumption which was shown to be potentially problematic in [67]). We show how to relax this assumption by assuming that the min-entropy rate of the QRNG output is only $r=0.99$, i.e. $k_{2}=r\cdot n_{2}$, and adjusting the output length of the extractor accordingly (in the quantum-proof product source model), see Figure 3. The code to perform the parameter calculation is given in Figure 6, but it could also easily be performed using the calculation module from_params (see documentation).
@@
(after sifting, error correction and parameter estimation).
@@
# To use Circulant, the seed length needs to be a prime with
# primitive root 2, we find the one larger than or equal to
# n_1 + 1 (see Def. 9):
n_adjusted = cryptomite.utils.next_na_set(n_1 + 1)
@@
k_2 = r * n_adjusted
# Compute the extractors output length (product source